OpenStudy (anonymous):

What is the product of 5the square root of u. (the square root of 3. + the square root of 147.)? Simplify if possible. 5the square root of 3 times u. + 7the square root of 3. 40the square root of 3 times u. 54the square root of 3 times u. 35uthe square root of 3.

OpenStudy (anonymous):

\[5\sqrt{u}(\sqrt{3}+\sqrt{147})\]?

OpenStudy (anonymous):

that is my guess anyway \[\sqrt{147}=\sqrt{49\times 3}=7\sqrt{3}\] so inside the parentheses is \(\sqrt{3}+7\sqrt{3}=8\sqrt{3}\) and \[5\sqrt{u}\times 8\sqrt{3}=40\sqrt{3u}\]